OpenStudy (anonymous):

(x + 11)(x - 11) is x^2 + 22

OpenStudy (anonymous):

thanks :)

OpenStudy (diyadiya):

x^-121

hero (hero):

x^2 - 121

OpenStudy (diyadiya):

it isn't 22

OpenStudy (anonymous):

how did you get 121

OpenStudy (diyadiya):

-11 * 11 =121

hero (hero):

difference of squares (a+b)(a-b) = a^2 - b^2

OpenStudy (anonymous):

ok i see

hero (hero):

x = a b = 11

hero (hero):

Mya, I know where your school is

hero (hero):

I'm coming to class